# Two-stage Estimation for Causal Inference Involving a Semi-continuous Exposure

### v2

@@ 2.2 Model Settings
with $\bar{\mathbf{Z}}_{2}=(1,\mathbf{Z}_{2}^{\prime})^{\prime}$ and $\boldsymbol{\alpha}_{2}=(\alpha_{20},\boldsymbol{\alpha}_{21}^{\prime},\boldsymbol{\alpha}_{22}^{\prime})^{\prime}$ where $\boldsymbol{\alpha}_{21}=(\alpha_{211},...,\alpha_{21k_{2}})^{\prime}$ and $\boldsymbol{\alpha}_{22}=(\alpha_{221},...,\alpha_{22k_{3}})^{\prime}$.
@@ 6.1 Description of the data
We adjust for baseline maternal sociodemographic characteristics, smoking and other substance use during pregnancy, reproductive history, gestational age at screening, and measures of the home environment and maternal cognition (listed in Table 7 in Appendix G.1). To address missingness, we generate 20 multiply imputed datasets using an imputation model including all variables; further details are given in Appendix G.2. Point estimates and standard errors are combined using Rubin’s rules under a missing-at-random assumption (Rubin, 1976; Little and Rubin, 2019).
@@ 6.2 Application and findings
Table 7 in Appendix G.1 summarises the fitted models for the continuous and binary exposure components, $E(D|A=1,\mathbf{X})$ and $P(A=1|\mathbf{X})$. Both models include all candidate covariates, following the recommendation of Brookhart et al. (2006) that including variables related to the outcome can improve precision of causal effect estimates.
@@ Acknowledgements
Appendix A. Identifying Assumptions
Assumption 5 (Stable Unit Treatment Value Assumption)
@@ Appendix A. Identifying Assumptions
Appendix B. Derivation of the Marginal Structural Model Parameters
@@
Appendix C. Proof of Balancing for Regression Adjustment with Two Propensity Scores
@@ C.2 Effect of Exposure at the Reference Dose versus No Exposure
Appendix D. Additional Materials on Statistical Inference
@@ D.2 Proof of Theorem 2
where $w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)={I\left(A_{i}=a\right)}/\left[{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)^{a}\left\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\right\}^{1-a}}\right].$ The first term in (D.2 Proof of Theorem 2) can be represented as
@@
Thus, we only need to show the second term in (D.2 Proof of Theorem 2) is equal to $\boldsymbol{0}$ under the following three cases.
@@
The second component in (D.2 Proof of Theorem 2) is
@@
Under (D.13), (D.2 Proof of Theorem 2) equals 0. Therefore, we conclude that $\hat{\gamma}_{21}$ is consistent for $\psi_{11}$ in the MSM (2.4) if the propensity score model is misspecified and the imputation model is correctly specified, under Assumptions 1–4 and based on a consistent $\hat{\gamma}_{11}$.
@@
The second component in (D.2 Proof of Theorem 2) can be re-expressed as
@@
Therefore, (D.2 Proof of Theorem 2) equals 0 in this case. In conclusion, $\hat{\gamma}_{21}$ is consistent for $\psi_{11}$ in the MSM (2.4) if the imputation model is misspecified and the propensity score model is correctly specified, under Assumptions 1-4 and based on a consistent $\hat{\gamma}_{11}$.
@@ D.4 Sandwich Variance Estimator and Wald Inference
Appendix E. Proof of the Bias Induced by Misspecified Propensity Scores in Regression
@@ E.1 Bias Caused by Misspecified Propensity Scores in Stage I
From (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
@@
From (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
@@
Then, we replace $\gamma_{10}$ in (E.1 Bias Caused by Misspecified Propensity Scores in Stage I)
@@
After replacing $\gamma_{10}$ in (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
@@
Similarly, after some algebra, from (E.1 Bias Caused by Misspecified Propensity Scores in Stage I) we finally get
@@
By replacing $\gamma_{12}$ in (E.1 Bias Caused by Misspecified Propensity Scores in Stage I), we get
@@ E.2 Bias Caused by Misspecified Propensity Scores in Stage II
Our next step is to take expectation with respect to $\mathbf{X}$. From (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) we get
@@
From (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) we get
@@
From (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) we get
@@
We replace $\gamma_{20}$ in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) and get:
@@
To solve for $\gamma_{22}$, we replace $\gamma_{20}$ in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) to get
@@
where $\delta_{1}=\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$ and $\delta_{2}=\text{cov}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$. Next, we substitute $\gamma_{22}$ in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II):
@@
Here we simplify some terms in (E.2 Bias Caused by Misspecified Propensity Scores in Stage II),
@@
where $\phi_{2j}(\mathbf{X})=\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}_{j}\bigr\}$ for j=1,2,,kj=1,2,...,k. Then (E.2 Bias Caused by Misspecified Propensity Scores in Stage II) can be expressed as
@@
Appendix F. Additional Materials on Simulation Study
@@ F.2 Additional Simulation Results
Table Table 1: Empirical results for estimators of causal effects with $\pi^{*}=0.25$ under six methods of analysis based on correctly specified models.
@@
Table Table 2: Empirical results for estimators of causal effects with $\pi^{*}=0.25$ for two-propensity-score regression adjustment and two-stage analysis under model misspecification.
@@
Table Table 3: Empirical results for estimators of effects for both parts with $\pi^{*}=0.25$ for applying PS regression adjustment in Stage I and AIPW in Stage II under model misspecification.
@@
Table Table 4: Empirical results for estimators of causal effects with $\pi^{*}=0.75$ under six methods of analysis based on correctly specified models.
@@
Table Table 5: Empirical results for estimators of causal effects with $\pi^{*}=0.75$ for two-propensity-score regression adjustment and two-stage analysis under model misspecification.
@@
Table Table 6: Empirical results for estimators of effects for both parts with $\pi^{*}=0.75$ for applying PS regression adjustment in Stage I and AIPW in Stage II under model misspecification.
@@
Appendix G. Additional Information about the Application
@@ G.1 Effects of Prenatal Covariates on PAE
Table Table 7: Estimated effects of prenatal covariates on the continuous and binary exposure components
@@ G.3 Details of Method Implementation in Data Application
G.3.1 Conventional Covariate Regression Adjustment.
@@
G.3.2 Two-Propensity-Score Regression Adjustment.

### v3

@@ 2.2 Model Settings
with $\bar{\mathbf{Z}}_{2}=(1,\mathbf{Z}_{2}^{\prime})^{\prime}$ and $\boldsymbol{\alpha}_{2}=(\alpha_{20},\boldsymbol{\alpha}_{21}^{\prime},\boldsymbol{\alpha}_{22}^{\prime})^{\prime}$ where $\boldsymbol{\alpha}_{21}=(\alpha_{211},...,\alpha_{21k_{2}})^{\prime}$ and $\boldsymbol{\alpha}_{22}=(\alpha_{221},...,\alpha_{22k_{3}})^{\prime}$. appendix
@@ 6.1 Description of the data
We adjust for baseline maternal sociodemographic characteristics, smoking and other substance use during pregnancy, reproductive history, gestational age at screening, and measures of the home environment and maternal cognition (listed in Appendix Table 7 in Appendix G.1). To address missingness, we generate 20 multiply imputed datasets using an imputation model including all variables; further details are given in Appendix G.2. Point estimates and standard errors are combined using Rubin’s rules under a missing-at-random assumption (Rubin, 1976; Little and Rubin, 2019).
@@ 6.2 Application and findings
Appendix Table 7 summarises the fitted models for the continuous and binary exposure components, $E(D|A=1,\mathbf{X})$ and $P(A=1|\mathbf{X})$. Both models include all candidate covariates, following the recommendation of Brookhart et al. (2006) that including variables related to the outcome can improve precision of causal effect estimates.
@@ Acknowledgements
Appendix A Identifying Assumptions
Stable Unit Treatment Value Assumption
@@ Appendix A Identifying Assumptions
Appendix B Derivation of the Marginal Structural Model Parameters
@@
Appendix C Proof of Balancing for Regression Adjustment with Two Propensity Scores
@@ C.2 Effect of Exposure at the Reference Dose versus No Exposure
Appendix D Additional Materials on Statistical Inference
@@ D.2 Proof of Theorem
where $w_{i}\left(a;\boldsymbol{\alpha}_{2}\right)={I\left(A_{i}=a\right)}/\left[{S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)^{a}\left\{1-S_{2}\left(\mathbf{Z}_{i2};\boldsymbol{\alpha}_{2}\right)\right\}^{1-a}}\right].$ The first term in (D.2) can be represented as
@@
Thus, we only need to show the second term in (D.2) is equal to $\boldsymbol{0}$ under the following three cases.
@@
The second component in (D.2) is
@@
Under (D.13), (D.2) equals 0. Therefore, we conclude that $\hat{\gamma}_{21}$ is consistent for $\psi_{11}$ in the MSM (2.4) if the propensity score model is misspecified and the imputation model is correctly specified, under Assumptions 1–4 and based on a consistent $\hat{\gamma}_{11}$.
@@
The second component in (D.2) can be re-expressed as
@@
Therefore, (D.2) equals 0 in this case. In conclusion, $\hat{\gamma}_{21}$ is consistent for $\psi_{11}$ in the MSM (2.4) if the imputation model is misspecified and the propensity score model is correctly specified, under Assumptions 1-4 and based on a consistent $\hat{\gamma}_{11}$.
@@ D.4 Sandwich Variance Estimator and Wald Inference
Appendix E Proof of the Bias Induced by Misspecified Propensity Scores in Regression
@@ E.1 Bias Caused by Misspecified Propensity Scores in Stage I
From (E.1), we get
@@
From (E.1), we get
@@
Then, we replace $\gamma_{10}$ in (E.1)
@@
After replacing $\gamma_{10}$ in (E.1), we get
@@
Similarly, after some algebra, from (E.1) we finally get
@@
By replacing $\gamma_{12}$ in (E.1), we get
@@ E.2 Bias Caused by Misspecified Propensity Scores in Stage II
Our next step is to take expectation with respect to $\mathbf{X}$. From (E.2) we get
@@
From (E.2) we get
@@
From (E.2) we get
@@
We replace $\gamma_{20}$ in (E.2) and get:
@@
To solve for $\gamma_{22}$, we replace $\gamma_{20}$ in (E.2) to get
@@
where $\delta_{1}=\text{cov}\bigl\{E(A|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$ and $\delta_{2}=\text{cov}\bigl\{E(AD|\mathbf{X}),\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2})\bigr\}$. Next, we substitute $\gamma_{22}$ in (E.2):
@@
Here we simplify some terms in (E.2),
@@
where $\phi_{2j}(\mathbf{X})=\text{cov}\bigl\{\tilde{S}_{2}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{2}),\mathbf{X}_{j}\bigr\}$ for j=1,2,,kj=1,2,...,k. Then (E.2) can be expressed as
@@
Appendix F Additional Materials on Simulation Study
@@ F.2 Additional Simulation Results
Appendix Table 1: Empirical results for estimators of causal effects with $\pi^{*}=0.25$ under six methods of analysis based on correctly specified models.
@@
Appendix Table 2: Empirical results for estimators of causal effects with $\pi^{*}=0.25$ for two-propensity-score regression adjustment and two-stage analysis under model misspecification.
@@
Appendix Table 3: Empirical results for estimators of effects for both parts with $\pi^{*}=0.25$ for applying PS regression adjustment in Stage I and AIPW in Stage II under model misspecification.
@@
Appendix Table 4: Empirical results for estimators of causal effects with $\pi^{*}=0.75$ under six methods of analysis based on correctly specified models.
@@
Appendix Table 5: Empirical results for estimators of causal effects with $\pi^{*}=0.75$ for two-propensity-score regression adjustment and two-stage analysis under model misspecification.
@@
Appendix Table 6: Empirical results for estimators of effects for both parts with $\pi^{*}=0.75$ for applying PS regression adjustment in Stage I and AIPW in Stage II under model misspecification.
@@
Appendix G Additional Information about the Application
@@ G.1 Effects of Prenatal Covariates on PAE
Appendix Table 7: Estimated effects of prenatal covariates on the continuous and binary exposure components
@@ G.3 Details of Method Implementation in Data Application
G.3.1 Conventional Covariate Regression Adjustment
@@
G.3.2 Two-Propensity-Score Regression Adjustment